OpenStudy (anonymous):

What is an equation in slope-intercept form for the line that passes through the points (1, –3) and (3, 1)

OpenStudy (anonymous):

y=mx+b m=y(2)-Y(1)/x(2)-x(1)

OpenStudy (anonymous):

The slope of the line is {1-(-3)}/(3-1)=2. Therefore y=2x +c. substitute (3,1) in this as the point lies on the line. 1=6+c. c=-5 eqn required is y=2x-5

OpenStudy (anonymous):

the 1's and 2's are subscripts

OpenStudy (cwrw238):

slope = difference in y values / difference in x values == (1-(-3) / 3 -1) = 4/2 = 2 now use general form y-y1 = m(x-x1) where m=slope and (x1,y1) = (3,1)